OpenStudy (jkristia):

I'm having a problem finding the limit when the function includes a trig function. E.g. page 74, #20 a, lim (x -> 0) of 'sin x / (3sqrt(x))' or #20 b, lim (x -> 0) of 'sin^2 x / x' Any hints on how to tackle these ?

OpenStudy (espex):

You attack it in the same manner you would if it was x^2 or any other function of x. As you apply the limit you will end up taking the sin(0). Use a substitution variable if that helps during the rationalization of the numerator, then just place the trig function back afterwards.

OpenStudy (jkristia):

eSpecX, I'm not sure I understand your explanation, but I found some helpful videos on youtube taking limits of trig functions. The goal is to always get the equation into the form of either sin x / x = 1 or (1 - cos x) / x = 0. In the case of \[\sin x / 3\sqrt{x}\] I should multiply by \[x / x\] to get\[(x / 3)(\sin x / x) = (x / 3)(1) = 0/3 = 0\] I think that is how it should be solved.